### 24 Sep 2025 v2

# When Your Boss Is an AI Bot: Exploring Opportunities and Risks of Manager Clone Agents in the Future Workplace

Abstract.
As Generative AI (GenAI) becomes increasingly embedded in the workplace, managers are beginning to create Manager Clone Agents—AI-powered digital surrogates that are trained on their work communications and decision patterns to perform managerial tasks on their behalf. To investigate this emerging phenomenon, we conducted six design fiction workshops (n = 23) with managers and workers, in which participants co-created speculative scenarios and discussed how Manager Clone Agents might transform collaborative work. We identified four potential roles that participants envisioned for Manager Clone Agents: proxy presence, informational conveyor belt, productivity engine, and leadership amplifier, while highlighting concerns spanning individual, interpersonal, and organizational levels. We provide design recommendations envisioned by both parties for integrating Manager Clone Agents responsibly into the future workplace, emphasizing the need to prioritize workers’ perspectives, strengthen interpersonal bonds, and enable flexible clone configuration.
Manager Clone Agents, AI Agent, Future of Work, Generative AI, Workplace, Human–AI Collaboration
conference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYccs: Human-centered computing Computer supported cooperative workccs: Human-centered computing Empirical studies in HCI
1. Introduction
Recent advances in AI have given rise to the phenomenon of managers “cloning themselves” to delegate routine and strategic work such as attending meetings, exchanging information with employees and stakeholders, and assisting in decisions (Forsdick, 2025; Raconteur, 2023; Gani, 2025). Unlike algorithmic systems that primarily optimize decisions or workflows (Lee et al., 2015), Manager Clone Agents replicate a manager’s symbolic, social, and communicative presence. They act as digital clones (Liang, 2025; Lee et al., 2023), mimicking a manager’s appearance, behaviors, and communication style, and can also evolve into functional delegates (Cheng et al., 2025; The Journey Platform, 2025) that perceive context, reason through tasks, and execute complex actions without intervention.
In practice, such systems are already emerging in heterogeneous forms, from virtual avatars that project an executive’s presence to lightweight assistants embedded in workflows that handle approvals or reports. These designs differ in fidelity, autonomy, but collectively signal a rapidly expanding design space. High-profile executives such as Nvidia CEO Jensen Huang (NVIDIA, 2022) and LinkedIn co-founder Reid Hoffman (HeyGen, 2024) have released their AI avatars in public, while companies like Zoom (Peters, 2025) and Otter.ai (Boyle, 2025) pilot digital stand-ins for meetings and briefings. Surveys further suggest that adoption is cascading beyond the C-suite into middle management (MASL WORLD, 2025; Langley, 2025), aided by falling implementation costs and expectations of measurable productivity gains.
Yet alongside this momentum lie pressing concerns. Because managers shape decision-making, resource distribution, and workplace culture (Mintzberg et al., 1973), delegating aspects of their authority, whether symbolically or functionally, to AI agents raises difficult questions of trust, fairness, and accountability. Workers worry about displacement and the erosion of authentic communication (Bonos and Abril, 2025). These tensions make it urgent to study how such agents alter leadership, collaboration, and organizational dynamics, and to identify pathways for their responsible design and adoption.
Existing HCI studies have begun to examine how AI Clones, personalized AI mimics that approximate individuals given limited training data (Liang, 2025), coexist with their human originals, often focusing on questions of identity, self-perception and other interpersonal relationships (Hwang et al., 2024; Huang et al., 2025; Wang et al., 2025; Lee et al., 2023). Building on this work, we use the term Manager Clone Agents to refer broadly to a family of emerging systems, ranging from symbolic clones that replicate presence to functional delegates that automate managerial tasks. Positioned at the intersection of authority and collaboration, they combine task-oriented delegation with the symbolic and communicative authority of human managers. This dual capacity makes them distinct from prior managerial technologies (Lee et al., 2015; Wolf and Stock-Homburg, 2025; Sowa et al., 2020), but we know little about how their human-like qualities will reshape teamwork within organizations.
Therefore, in this paper, we ask:
  • RQ1: What opportunities and supportive roles do managers and workers envision for Manager Clone Agents in the future workplace?
  • RQ2: What challenges and risks do managers and workers associate with the use of Manager Clone Agents in the future workplace?
  • RQ3: How do managers and workers envision the future design of Manager Clone Agents to address these opportunities and risks?
To explore these questions, we conducted a series of design fiction workshops (Lyckvi et al., 2018; Muller et al., 2020) with managers and workers, asking participants to imagine organizational futures where Manager Clone Agents were widely adopted. Design fiction, a method widely used in HCI, was chosen because Manager Clone Agents remain speculative technologies in this moment for most managers and workers: it allows participants to engage with fictional yet plausible scenarios to articulate hopes, conflicts, and concerns without being constrained by today’s technical limits. In this paper, we define “managers” as organizational leaders at multiple levels, including executives, directors, and middle managers, and “workers” as their employees and direct reports who experience the effects of managerial authority. In the workshops, both groups created scenarios of Manager Clone Agents use and then co-reflected on the opportunities, tensions, and uncertainties these agents might generate in organizational life.
Our findings revealed a dual character to Manager Clone Agents. On one hand, participants envisioned supportive roles, acting as proxy presences in meetings, conveying information across organizational layers, automating routine tasks, and amplifying managerial guidance (RQ1). On the other hand, these same roles raised risks across multiple levels: at the individual level, from different positions both managers and workers experienced shared anxieties about accountability and irreplaceability; at the interpersonal level, both sides pointed to fragile trust, blurred authenticity, and reduced informal contact; and at the organizational level, participants worried that efficiency gains could flatten hierarchies and weaken belonging (RQ2). Participants therefore stressed that responsible design must foreground workers’ voices, preserve communication and social cohesion, and allow trust to develop gradually through clear boundaries on what Manager Clone Agents can and cannot do (RQ3).
This study makes three key contributions to HCI and CSCW community:
  • First, we present empirical vision, shaped together by managers and workers of potential roles that emerging Manager Clone Agents could undertake and surface the opportunities and risks these roles create across individual, interpersonal, and organizational levels in future workplaces.
  • Second, we distinguish Manager Clone Agents from prior algorithmic management systems by showing how their personalization and symbolic authority recast debates on authenticity and automation, while reframing both managerial work and worker-manager relations as relational and organizational challenges of AI automation.
  • Finally, we surface design directions for developing Manager Clone Agents that build trust and evolve collaborations within diverse team and organizational contexts.
2. Background: Emerging Manager Clone Agents in Workplaces
Recent advances in artificial intelligence have enabled the rise of Manager Clone Agents. Executives are beginning to use these tools to manage heavy schedules, from financial briefings to routine worker inquiries. As early as 2021, Nvidia CEO Jensen Huang released an AI avatar powered by Omniverse Avatar (NVIDIA, 2022). Since then, startups such as Personal AI [^1] https://www.personal.ai/ , Delphi [^2] https://www.delphi.ai/ , and Tavus [^3] https://www.tavus.io/ have built platforms that train avatars on leaders’ speeches and meetings to create chatbots, voice agents, and video avatars. Investment has followed quickly: Delphi raised \$16M in June 2025 (Ladjevardian, 2025), HeyGen \$60M (Team, 2025), and ElevenLabs \$180M at a \$3.3B valuation (Mehta and Lunden, 2025), while Synthesia reached \$2.1B (Lunden, 2025). High-profile leaders such as Linkedin co-founder Reid Hoffman, Zoom CEO Eric Yuan, and Khosla Ventures managing director Keith Rabois have also begun experimenting with these systems (HeyGen, 2024; Patel, 2024; Morrone, 2025), using them for financial updates (Bort, 2025; Peters, 2025), meeting attendance (Boyle, 2025), or routine staff communication (Personal AI, 2025; Ladjevardian et al., 2025; Langley, 2025; Cuthbertson, 2023). Adoption of AI agents in workplaces is also accelerating. A PwC survey of 308 U.S. managers in April 2025 reported that 79% of companies already use AI agents, and 88% plan to expand budgets within a year (PwC, 2025). Falling costs are lowering entry barriers (Fifer and Feagans, 2025), positioning Manager Clone Agents to spread from executives to middle and line managers in the near future (DigitalDefynd, 2025). Yet concerns remain. Because managerial roles shape decisions, resources, and culture, delegating them to AI raises questions of trust, accountability, and fairness (Sowa et al., 2020). Scholars warn that these systems may reinforce executive detachment (Morrone, 2025), while workers fear job loss and erosion of authentic communication (Writer, 2025). These tensions highlight the need to better understand how Manager Clone Agents affect collaboration and organizational life.
3. Related Work
First, we review scholarship on managerial roles in the workplace, which highlights the complex mix of decision-making, coordination, and relational care that define managerial labor (see subsection 3.1). Second, we discuss HCI and CSCW work that examines the role of technologies in supporting managerial work, from algorithmic management to collaborative robots and AIs (see subsection 3.2). Third, we examine early work on Clone Agents studies, focusing on their potential to extend human presence while raising new concerns around authenticity (see subsection 3.3).
3.1. Managerial Roles in the Workplace
Managers occupy a distinctive position in organizational life. Unlike most employees, they are responsible for abstract decision-making, conflict coordination, and emotional motivation, which make their tasks qualitatively different from those of their subordinates (Dler and Tawfeq, 2021). Classic organizational theory frames managerial work as a constellation of intertwined roles: interpersonal (figurehead, leader, liaison), informational (monitor, disseminator, spokesperson), and decisional (entrepreneur, disturbance handler, resource allocator, negotiator) (Mintzberg et al., 1973). Yet these responsibilities accumulate into heavy burdens. Managers already face fragmented routines of meetings, emails, and constant communication that limit time for reflection and strategy (Stray and Moe, 2020; Dler and Tawfeq, 2021). Recent shifts in digital technologies that multiply channels and complicate coordination (Andersone et al., 2022), along with the demands of remote and hybrid work (Morrison-Smith and Ruiz, 2020; Choudhury et al., 2023) have further amplified these strains.
Beyond the structural demands of the role, effective management depends on the individual ways managers enact their responsibilities. Leadership research highlights how different decision-making styles (e.g., autocratic, democratic) produce divergent effects on team performance, satisfaction, and trust (Yukl and Fleet, 1992). Some managers also construct personal narratives that align organizational goals, thereby legitimizing their authority (Rostron, 2022). These perspectives underscore that effectiveness in management hinges less on a fixed formula and more on how managers’ personal identities and the relational dynamics they cultivate. Organizational efforts to ease workload, such as appointing chiefs of staff, executive assistants, or project management offices, reinforce this point: the success of such delegation hinges on trust, specifically on whether the decision-making comes from trusted internal leaders or less familiar external consultants (Meagher and Wait, 2020; Tamkin, 2014).
This body of work shows that managers must simultaneously juggle strategic, operational and interpersonal duties, while also enacting identities that sustain their authority. The sophistication and personalization of management, and the role it plays in team dynamic, call for more nuanced and effective support. In this paper, we explore Manager Clone Agents as a novel possibility: digital delegate trained to mimic a manager’s communication style, decision-making patterns, and even personal presence.
3.2. Technologies that Support Managerial Roles
Scholars in HCI and CSCW have long examined how algorithmic systems, robots and other forms of AIs assume partial managerial roles, reshaping workflows and authority structures. One prominent line of research is algorithmic management (Lee et al., 2015; Lee and Baykal, 2017; Lee et al., 2019a; Green and Chen, 2019; Cao et al., 2021), where decision-making roles are automated through software and real-time data analytics. It has underscored how algorithmic oversight reconfigures worker-manager relations, offering both tighter control and diminished autonomy for human supervisors. Another related stream focuses on how collaborative robots (cobots) (Cheon et al., 2022) reshaped roles and trust, and that an ideal image of a robot leader needs to be tailored to individuals and set appropriate expectation management (Wolf and Stock-Homburg, 2025).
AI is no longer merely an assistive tool that teams use to accomplish tasks; it is increasingly becoming an active part of the team, functioning as an agent that works alongside people in their everyday collaboration (Ofem, 2025; Dennis et al., 2023). At the executive level, recent scholars have begun theorizing about the Chief AI Officer (CAIO) role as a strategic necessity (Schmitt, 2024). Research in Human–AI Teaming (HAT) further shows that when AIs join teams, the initial expectation they received in terms of communication capabilities, human-like behaviors and performance shape how trust and power are formed and developed among the team (Duan et al., 2024; Zhang et al., 2021; Flathmann et al., 2024). Building on this, Sadeghian et al. highlights how the interaction patterns between workers and AI systems directly influence workers’ sense of satisfaction and the meaning they derive from their roles (Sadeghian et al., 2024). These studies establish that technologies, whether algorithms, collaborative robots, or AI leadership frameworks, do more than optimize processes: they reshape authority, identity, and social dynamics within organizations.
At the same time, the introduction of collaborative and managerial technologies provokes persistent concerns among workers about automation, control, and trust. Grudin’s well-known observation in 1988 has already warned that groupware adoption often fails when those responsible for entering data do not benefit directly from its use, or when organizational cultures prioritize competition over sharing (Grudin, 1988). Recent case studies extend these concerns to AI-based systems, showing that successful adoption depends on managers addressing both cognitive trust (clear understanding of how AI works and what its limitations are) (Ofem, 2025) and emotional trust (a sense of safety, fairness, and empathy in its use) (Vuori et al., 2025; Lee, 2018; Rosenblat and Stark, 2016). In particular, the negative impacts of algorithmic management often arise from systems that favor managerial efficiency over an inclusive design that centers on worker well-being (Zhang et al., 2022; Spektor et al., 2023). In many cases, the success or failure of management technologies hinges not only on technical accuracy but also on how well they align with cultural values of fairness, agency, and inclusiveness.
These streams of research reveal that technological interventions in management often boost efficiency and control but risk rigidity, impersonality, and worker resistance. They also prioritize procedural and physical aspects over relational and cultural dynamics. As collaborative technologies develop, leadership requires flexibility, trust-building, and shared meaning. Manager Clone Agents, unlike rigid algorithms, offer personalized support by simulating real managers’ styles to mediate tasks, communication and collaboration.
3.3. Clone Agents in Workplace
With the rise of GenAI, AI agents have emerged as a distinct category of intelligent digital assistants, capable of autonomously performing tasks on behalf of users and mimicking aspects of human interaction (Park et al., 2023, 2024; Li et al., 2024). A growing stream of research explores their role in teamwork, where AI agents can support managerial functions ranging from scheduling (Cranshaw et al., 2017) to decision-making (Li et al., 2025). Systems such as Cobot-style workflow assistants (Sowa et al., 2020) illustrate how agents can reduce leaders’ follow-up burdens. Other studies highlight how conversational agents can facilitate group discussions, such as encourage equal participation (Do et al., 2022; Kim et al., 2020), and team cohesion (Benke et al., 2020). These developments signal the growing managerial relevance of AI agents in collaborative contexts.
Building on this trajectory, Clone Agents represent a special subtype of AI agents. Unlike generic digital assistants, Clone Agents are designed to replicate not only an individual’s decision logic (Truby and Brown, 2021) but also their voice, appearance, and conversational style (Dobre et al., 2022; Leong et al., 2024; Shamekhi et al., 2018). Scholars have begun to conceptualize these entities as AI clones—digital representations trained on personal data that can interact with others as if they were the real-world individual (Lee et al., 2023). Early HCI and CSCW systems hinted at this direction: Time-Travel Proxy enabled asynchronous participation through pre-recorded video (Tang et al., 2012), and Lee & Takayama showed how mobile remote presence let remote workers “live and work” with colleagues almost as if physically there (Lee and Takayama, 2011). More recent work extends these ideas: Dittos created embodied proxies for real-time meetings (Leong et al., 2024), and Cheng et al. studied how delegating conversational autonomy to agents affects multitasking in group interactions (Cheng et al., 2025).
Manager Clone Agents, unlike rigid algorithms, offer personalized support by simulating real managers’ styles to mediate tasks, communication and collaboration (Forsdick, 2025). Whereas digital twins (Vainionpää et al., 2022; Lv et al., 2023) primarily model physical or organizational assets, algorithmic management (Lee et al., 2015; Zhang et al., 2022) emphasizes efficiency and control, and avatars (Gao et al., 2020; Weidner et al., 2023) or cobots (Weiss et al., 2021; Wolf and Stock-Homburg, 2025) focus on embodiment and interaction, Clone Agents are distinctive in that they merge personalization with symbolic authority, functioning as proxies that extend a specific manager’s presence and decision-making into organizational life. This dual capacity makes them qualitatively different from prior technologies, raising new questions about authenticity, legitimacy, and relational trust in AI-mediated management.
In terms of identity and authenticity, a core concern is that Clone Agents can exploit or steal an individual’s identity (Huang et al., 2025; Wang et al., 2025). Some fear “doppelganger-phobia”: one’s digital clone might replace or outperform them, potentially altering relationships with colleagues, employers, or even loved ones (Lee et al., 2023). Trust also becomes vulnerable in environments where clones and deepfakes circulate as synthetic media erodes confidence in visual and auditory evidence (Johnson and Diakopoulos, 2021; Kaate et al., 2023; Hwang et al., 2024). Finally, privacy and consent emerge as pressing challenges, as digital likenesses may be generated or deployed without authorization (Umbach et al., 2024; Lee et al., 2023).
Clone Agents hold significant promise as a novel form of AI agent in the future workplace, offering new ways to extend presence, delegate routine managerial work, and maintain continuity in collaborative settings. Yet, their uniqueness also introduces profound risks for accountability, trust, and word culture. Within this landscape, Manager Clone Agents mark a critical turning point, concentrating these opportunities and risks in ways that directly challenge how authority and collaboration are organized in the future workplace.
4. Method
To investigate how Manager Clone Agents might reshape collaboration in the workplace, we conducted six two-hour design fiction workshops with 23 participants in total, comprising both managers and workers in each session. Each workshop brought together three to five participants who co-created speculative scenarios and reflected on the opportunities, conflicts, and risks that Manager Clone Agents might generate in organizational contexts. Design fiction was particularly well-suited to this study because Manager Clone Agents remain speculative technologies (Muller et al., 2020; Lyckvi et al., 2018); by situating participants in fictional, yet plausible scenarios, we were able to elicit relational concerns and organizational imaginaries that might not emerge through surveys or interviews focused only on current practices.
4.1. Participants
In this study, we define managers as individuals who hold formal authority to supervise, coordinate or make decisions affecting others’ work, and workers as those who carry out tasks under managerial oversight without primary supervisory responsibilities. We specifically recruited both managers and workers to capture perspectives from those exercising managerial authority as well as those subject to it.
We recruited 23 participants through professional networks, workplace mailing lists, and social media outreach. To ensure diversity, we used purposive sampling (Campbell et al., 2020) to include individuals from a range of industries (e.g., technology, design, education, media, healthcare, manufacturing, and government) and with varied levels of work experience (from less than one year to over a decade). Table 1 summarizes participants’ demographic details.
Table 1. Workshop Participants’ Demographic.
ID Sex Age Work Experience Industry Team
W1 Female 28 1–2 years Design Product Team
M1 Male 25 1–2 years Renewable Energy Risk Management Team
W2 Female 24 1–2 years Automotive Manufacturing Design Team
M2 Female 29 3–5 years Automotive Manufacturing Design Team
M3 Male 25 3–5 years Film and Television Production Technical Production Team
M4 Male 24 1–2 years Media Event Planning Team
W3 Female 24 0–1 year Internet Data Production Team
W4 Female 26 0–1 year Technology Research Team
M5 Female 30 1–2 years Education Product Team
M6 Male 28 3–5 years Internet Cross-Functional Team
W5 Male 25 0–1 year Technology and Education Research Team
W6 Female 26 0–1 year Internet Product Team
M7 Female 27 1–2 years Government Agencies Project Execution Team
W7 Male 28 3–5 years Digital Media Editorial Team
W8 Male 27 3–5 years Healthcare Design Team
M8 Female 30 3–5 years Professional Development Cross-functional Team
M9 Female 31 5–10 years Marketing Marketing Team
M10 Male 32 5–10 years Foundation Model Infrastructure Software QA Team
M11 Male 45 ¿10 years Information Technology Product Team
W9 Female 25 1–2 years Racing Equipment Product Team
M12 Male 39 5–10 years Design Communication Design Team
M13 Female 33 5–10 years Technology UX/UI Design Team
W10 Female 27 3–5 years Design UX/UI Design Team
Note: The “ID” prefix indicates participant’s role (M = Manager, W = Worker). Participants grouped under each section participated in the same Workshop.
Our sample included 13 managers (M1 to M13) and 10 workers (W1 to W10). Participants ranged in age from 24 to 45 years across sectors. Notably, some workers in our sample had relatively long professional experience, while some managers were comparatively early in their careers. This reflected the fact that managerial status was not strictly tied to years of employment but to current responsibilities: in some organizations, employees became team leads or middle managers within just a few years, while in others, individuals with longer tenure continued to work as direct contributors under managerial supervision. Role categorization in our study therefore followed participants’ self-identification based on their formal position and everyday duties, rather than professional tenure alone. During the workshops, participants were asked to fully inhabit the roles they currently held within their company or team, drawing on their lived experiences as managers or workers to co-create and reflect on speculative scenarios of imagined futures about Manager Clone Agents.
4.2. Design Fiction Workshops
Each design fiction workshop lasted approximately 120 minutes, following a structured sequence of activities designed to elicit both imaginative and critical reflections on Manager Clone Agents in future teams and organizations. To mitigate potential power asymmetries between managers and workers, we ensured that each participant had dedicated turns to speak during group activities, explicitly invited quieter participants to share their views. Structured facilitation further emphasized that both managerial and worker perspectives carried equal value, reducing the likelihood that managerial voices would dominate. IRB approval was obtained prior to the workshops; all participants provided informed consent, and each received \$30 gift card after participation.
4.2.1. Introductions and role grounding.
At the start of each session, participants introduced themselves by describing their primary responsibilities within their current organization, their experiences of teamwork, and most importantly, their experiences of managing others or being managed. To anchor discussion in lived realities, each participant was also asked to share three to five concrete examples from their own work where managerial presence or absence significantly shaped outcomes.
4.2.2. Scenario brainstorming.
Building on these examples, participants individually brainstormed how their cases might unfold differently if a manager had been replaced by a Manager Clone Agent. Managers were asked to adopt the perspective of a leader delegating authority to a Manager Clone Agent, while workers were asked to imagine themselves being directed or supported by such an agent. This activity helped situate speculative thinking in the specific dynamics of authority, oversight, and collaboration familiar to each role.
4.2.3. Design fiction writing.
Each participant then selected one scenario to develop in greater depth. They spent 15 minutes writing a short design fiction narrative describing how future workplace practices, relationships, and outcomes would unfold when Manager Clone Agents were involved. Participants were encouraged to build directly on the examples they had shared earlier, but were also free to expand creatively and develop entirely new directions. This balance ensured that narratives remained grounded in lived experience while allowing space for imaginative exploration of near-future possibilities. It is noted that these design fictions are based on the imaginaries from our participants, instead of existing Manager Clone Agents deployments.
4.2.4. Story sharing and role-play discussion.
After writing, participants took turns presenting their stories to the group. Following each presentation, other participants engaged in a role-play discussion by responding as if they were members of that scenario, either managers or workers based on their own roles, reflecting on how they would experience and react to the presence of a Manager Clone Agent. This process encouraged participants to inhabit perspectives beyond their own and to surface tensions between managerial and worker vantage points.
4.2.5. Final round-table discussion.
Once all individual stories and discussions were completed, the group reconvened for a round-table discussion. Building on the scenarios they had created, participants reflected on several guiding questions. First, they considered what kinds of work or duties Manager Clone Agents could reasonably take over on behalf of managers, and which responsibilities should remain human-led. Second, they debated the benefits and concerns of using clone agents for managers, and what implications these might have at the individual (for both workers and managers), team, and organizational levels. Third, participants discussed how Manager Clone Agents should be designed in the future, including what forms of support would be most useful and what boundaries should be preserved. Finally, they compared how managers and workers viewed these issues differently, highlighting the distinct interests and considerations that shaped each group’s stance. These round-table conversations allowed participants to move from individual stories to broader reflections, surfacing shared opportunities, points of contention, and open questions about the integration of Manager Clone Agents in future workplace.
5. Data Analysis
We analyzed the workshop materials using reflexive thematic analysis (Braun and Clarke, 2021, 2019). Our dataset included participants’ written design fiction stories, notes taken during scenario brainstorming, and transcripts from group discussions and round-table synthesis. All workshops were audio-recorded and transcribed verbatim. To preserve the richness of participants’ accounts, we treated both fictional narratives and reflective commentary as complementary forms of data: the stories captured imaginative possibilities, while the discussions revealed participants’ evaluations, concerns, and reasoning.
The analysis followed an iterative, multi-stage process. First, two researchers independently read through all transcripts and narratives to gain familiarity with the data and to note initial impressions. Second, we conducted open coding at the sentence and paragraph level, tagging both descriptive content (e.g., tasks automated, benefits cited, risks identified) and relational dynamics (e.g., authority, trust, recognition). Third, we collaboratively grouped codes into higher-level categories that captured recurring patterns across workshops. Throughout this process, we regularly revisited raw data to ensure that emerging categories remained grounded in participants’ accounts.
In the final stage, we synthesized categories into three overarching themes that structured our findings: (1) the supportive roles participants envisioned for Manager Clone Agents (see section 6), (2) the risks and tensions they anticipated at individual, interpersonal, and organizational levels (see section 7), and (3) the future design directions they proposed (see section 8). These themes reflected both convergence and divergence in perspectives across managers and workers, highlighting how Manager Clone Agents were imagined not only as technical systems but as relational actors within organizational life. These findings themes are also summarized in Figure 2 for reference.
To enhance trustworthiness, the research team engaged in regular debriefing sessions to challenge assumptions and refine interpretations. Disagreements were resolved through discussion, and multiple iterations of coding and theme development were used to balance consistency with openness to new insights. Following a traditional HCI qualitative analysis approach (e.g., (Xiao et al., 2025b; Fan et al., 2025)), we did not assess statistical reliability or validity; instead, our reflexive thematic analysis emphasized depth of interpretation, contextual grounding, and transparency in the analytic process.
Figure 1. Summary of Findings.
6. Findings: Empowering Teamwork: Four Potential Supportive Roles of the Manager Clone Agents (RQ1)
Our findings reveal that participants envisioned Manager Clone Agents to empower future teamwork across four distinct supportive roles. As a Proxy Presence, they were expected to mitigate the costs of managerial absence by sustaining responsiveness and symbolic availability when managers were busy and not available (see subsection 6.1). Acting as an Information Conveyor Belt, they were imagined to streamline communication flows, reduce distortion, and surface grassroots concerns that might otherwise be lost in hierarchies (see subsection 6.2). As a Productivity Engine, they were expected to expand managers’ capacity by automating routine tasks while also freeing cognitive bandwidth for more strategic work (see subsection 6.3). Finally, as an Amplifier of Leadership, they were envisioned to extend the manager’s coaching, feedback, and mentoring to a broader set of workers, compensating for the limits of managerial time and attention and magnifying guidance that would otherwise remain scarce (see subsection 6.4).
6.1. A Proxy Presence: Always Present, Always on Call
Participants consistently described Manager Clone Agents as forms of proxy presence that could stand in for leaders when physical attendance was difficult or impossible. They envisioned these agents not as full replacements, but as a way to project presence, maintain communication, and preserve authority across different situations. This proxy presence was articulated in three ways: first, as a bridge during crises when managers could not be present; second, as a tool for more flexible and asynchronous participation in everyday work; and third, as a symbolic extension of managerial authority that differed in meaning from delegating to human assistants.
Many managers and worker participants in our study emphasized that the burdensome nature of leadership responsibilities often makes physical presence difficult. M6 described how managers like him are constantly on the move: “meeting clients, traveling across regions, and balancing competing responsibilities.” He added that this constant mobility left him mentally and physically exhausted, as he was expected to maintain visibility and responsiveness while being pulled across multiple sites of work. Other managers echoed this sentiment, noting that such demands delayed responses and reduced availability. Several workers, including W4, W6, and W8, also described feeling hesitant to approach managers who were perpetually “elsewhere”, worrying that asking for help would add to their burden.
In this context, participants saw Manager Clone Agents as a way to reduce the social costs of absence by providing a form of presence when managers could not attend in person. For example, many worker participants think, rather than leaving a meeting empty or canceled, the agent could appear with a simulated voice and avatar, signaling the manager’s involvement in a more respectful way than a brief apology about the disappearance.
M9, a marketing manager with 5–10 years of experience, illustrated this possibility in the speculative design story she wrote. In her scenario, a family emergency coincided with an important cross-timezone stakeholder meeting. Rather than canceling, she imagined sending her avatar, having pre-uploaded key talking points so that “the discussion could still move forward even if I wasn’t there in person.”
Beyond crisis situations, participants also saw Manager Clone Agents as tools for more flexible participation. By mediating communication asynchronously, they allowed both managers and workers to engage at their own pace rather than in real time. W8 explained: “With the Manager Clone Agent standing in, it’s the same as leaving a voice message. I can just listen whenever I have time, until it’s the best time for me.”
Additionally, compared with interacting directly with the real manager, some worker participants described feeling less pressure to give undivided attention when meeting with clone agents. As W8 put it, “If they’re talking about results, I’ll just leave it on in the background, like a Zoom meeting with camera and mic off.” This comment highlighted how the mediated presence of a Manager Clone Agent shifted expectations: workers could still access managerial input without the same demand for immediate, face-to-face attentiveness.
Participants also noted that the symbolic presence of a Manager Clone Agent carries different weight than sending an assistant or proxy human. As M12 put it, “Even though it’s a clone, it’s your leader’s clone, not another person who might distort information.” Here, the symbolic value lies in the continuity of identity: the clone embodies the manager’s voice and authority, even if mediated, whereas an assistant is seen as a separate actor. M10 believed that, for broadcast-style meetings or official statements, this symbolic link made the clone agent feel more authoritative and representative than delegating to other staff. At the same time, participants emphasized that symbolic presence has limits. As M4 noted, the manager’s personal appearance still matters in high-stakes contexts where trust cannot be fully substituted by a clone.
6.2. An Information Conveyor Belt: Filtering Noise and Freeing Voices
Participants repeatedly highlighted the communicative role of managers as conduits of information—“passing directives downward and carrying feedback upward, particularly for mid-level managers” as M6 described. They envisioned Manager Clone Agents as impartial conveyors that could transmit policies and feedback more clearly, reduce cognitive misalignment, and create safer channels for workers’ voices to be heard. Beyond individual exchanges, participants also imagined these agents improving large-scale information gathering, amplifying grassroots concerns that might otherwise be lost or filtered.
In discussing impartial transmission, M10 noted that messages relayed through human intermediaries were “less likely to be impartial”, often shaped by bias, incomplete understanding, or social discomfort. Building on this, workers like W3 and W4, along with managers such as M12, imagined Manager Clone Agents as more neutral carriers. In their stories, agents could smooth over gaps, keep meaning intact, and reduce the confusion that often arises when policies and feedback travel through layers of hierarchy.
Participants also reflected on how Manager Clone Agents might reduce the friction caused by cognitive misalignment between managers and workers. W4 recounted the difficulty of explaining her visa situation to a manager unfamiliar with international work rules: “Earlier I discussed with my manager how I could extend my OPT. Since he is not an international worker, I needed to provide him with some context and continue to tell him what my timeline looks like, and he has to continue to report to his manager and human resources.” In her view, the problem was not unwillingness but the sheer effort required to keep restating context across different levels of hierarchy. A Manager Clone Agent, participants suggested, could retain and transmit such contextual knowledge on behalf of the worker, ensuring that details are not lost each time the message is passed upward, while also sparing workers from the exhausting cycle of repeated clarification.
Another recurring theme was the potential to free voices that might otherwise be silenced by hierarchy. M10 observed that “in the workplace, workers may have some requests and not dare say them directly.” W9 in the same workshop echoed this, imagining Clone Agents as mediators that could create safer channels for raising concerns. In one speculative story, M2 described recruitment scenarios where applicants could ask sensitive questions more comfortably to an agent than to a manager, since the agent could respond with patience and without judgment, and applicants would not feel embarrassed or disadvantaged for raising such issues.
Finally, participants described how Manager Clone Agents might improve large-scale information collection. M12 imagined an agent preparing for a general assembly by first meeting with smaller groups to surface pressing issues: “To prepare for a general assembly, an Agent could meet with smaller groups (departments, floors, etc.) in order to figure out what pressing issues should be addressed.” W10, in the same workshop, agreed that this would make workers feel heard, contrasting it with surveys: “If you have the Manager Clone Agent, you will feel like someone is listening to you, and they will do something about it.”
6.3. A Productivity Engine: Automating Tasks and Augmenting Intelligence
Participants imagined Manager Clone Agents as productivity engines that could expand managerial capacity in two complementary ways: by automating repetitive, rule-based tasks and by extending cognitive functions that support decision-making. These dual capacities suggested a vision of leadership freed from low-level burdens and better equipped to focus on strategy, growth, and human relationships.
Many participants described how Manager Clone Agents could absorb what M7 called “high-frequency, low-risk, and rule-based” processes that otherwise drained managerial time and attention. Routine approvals, such as “approving invoices” (M4), “verifying expense claims” (M7), “processing leave” (W1, M6), or “signing off on weekly timesheets” (W6), were repeatedly mentioned as tasks that could be offloaded.
As M7 explained, these workflows were “essentially repetitive approvals that follow preset rules; having an agent check and approve them would save enormous effort without shifting responsibility.” Others emphasized that such automation could bring flexibility, with W1 noting that conflicts or sick leave could be handled more smoothly if agents were able to reschedule meetings or process minor requests.
In more complex contexts, participants even imagined agents carrying through contract reviews or resource allocations once decisions had already been settled. M7, for instance, told a story where “Party B mailed back the stamped contract (which had already been revised several times). They hope we approve and stamp it promptly and release the funds. At this point, the AI agent, based on the previously discussed details […] then proceeds with the stamping process, seeks finance approval, and completes the disbursement procedures.” She estimated that such automations could offload up to “10–15% of the workload.” Some, like M10, went further, imagining multi-agent setups where a Manager Clone Agent orchestrated specialized sub-agents for scheduling, finance, or reporting, thereby scaling efficiency beyond what one manager could achieve.
Participants further emphasized that agents might enhance fairness and accuracy in managerial decisions. M6 believed that by maintaining records and reducing reliance on subjective memory, agents could bring greater impartiality to areas such as task allocation or performance evaluation. M7 imagined them “reducing errors caused by fatigue, standardizing compliance checks,” while M3 suggested they could support “fairer, data-driven judgments.” At the same time, there are some cautions. M1 and W3 worried that delegating too much discretion could backfire if managers later overturned an agent’s decision, eroding credibility. For most of the participants, the promise of cognitive support had to be carefully balanced against the risk of eroding managerial authority.
6.4. An Amplifier of Leadership: Scaling Support, Enabling Growth
In our workshops, participants envisioned how Manager Clone Agents might amplify a manager’s impact by scaling day-to-day guidance to multiple workers at once, widening access to the manager’s attention, and creating space for workers to develop independent judgment. They emphasized that amplification holds potential for helping teams mature, provided that boundaries, escalation paths, and norms against over-reliance are made explicit.
Participants described Manager Clone Agents as useful for coordinating work and giving timely guidance so that tasks do not stall while waiting for a meeting. In W8’s story, the agent greeted him on Monday with a ranked task list, answered common questions, and, when reaching its limits, pinged the manager while suggesting productive interim tasks. W5 similarly imagined an agent from his supervisor accelerating turnaround times, while M1 described lightweight nudges such as a Slack bot that periodically reminds workers of handoffs or check-ins.
Amplification also meant making managers more accessible to workers. W3 explained that a Manager Clone Agent could capture feedback right after a meeting that might otherwise be lost if the real manager was traveling or tied up in back-to-back calls. W4 added that reaching an agent felt more approachable and responsive than sending what she called a “cold email.”
Managers stressed that this kind of amplification works only when agents enforce role boundaries and escalate issues at the right thresholds. M8 envisioned setting clear pre-approved decision bands in her story, noting that “My team can ask my agent if I am okay with certain decisions or then alert me if my authority or insight is needed for a potential project decision. My team would view this as potentially having greater access to me and the resources to get the job done. They would be able to feel more empowered to make decisions and could run thoughts by me more quickly.” At the same time, W10 cautioned that “There’s a difference between giving someone feedback and spoon-feeding them, the latter would create a reliance on constantly having someone to bounce ideas off 24-7.” If the agent is always available, juniors may defer decisions they should learn to own.
Finally, participants described softer forms of amplification, such as carrying culture and care. M5 imagined a Manager Clone Agent that “facilitates focus sessions and mirrors a mentor’s way of thinking on a mentee’s slides or writing”, thereby “passing on both ideas and spirit”. M10 shared how an automated birthday note from a system unexpectedly rekindled a relationship with a former teammate; others saw similar low-stakes, high-touch gestures as well-suited for Manager Clone Agents to maintain warmth and connection across large organizations.
7. Findings: Navigating Risks: Individual, Interpersonal, and Organizational Impacts (RQ2)
While the previous section highlighted how participants envisioned opportunities for extending managerial reach, here we turn to the risks that, in our participants’ imaginaries, could emerge if Manager Clone Agents were poorly integrated into everyday practices. These risks spanned individual, interpersonal, and organizational levels. At the individual level (see subsection 7.1), both managers and workers grappled with shared vulnerabilities around accountability, skill erosion, and replaceability, though their stakes diverged: managers feared threats to authority and identity, while workers worried about blocked career pathways and recognition. At the interpersonal level (see subsection 7.2), trust initially transferred from human relationships to Manager Clone Agent-mediated exchanges, but fragile authenticity, weakened affective cues, and reduced informal contact strained workplace bonds. At the organizational level (see subsection 7.3), efficiency gains were anticipated to flatten hierarchies and accelerate workflows, yet participants feared this would hollow out belonging and weaken culture.
7.1. Individual Risks: Divided Titles, Shared Vulnerability
Across discussions, managers and workers found themselves bound by similar anxieties yet separated by divergent stakes. Both recognized that Manager Clone Agents unsettle the rules of responsibility, reshape opportunities for growth, and test what it means to remain irreplaceable for human employees. But while managers viewed agents as a challenge to their authority and professional identity, workers saw them as blocking pathways to recognition and advancement. These overlapping perspectives reveal how Manager Clone Agents use exposes both shared vulnerabilities and role-specific divides.
Participants often returned to the question of accountability, disagreeing on who carries the blame when a Manager Clone Agent speaks or acts. Concerns centered on errors: agents misinterpreting intent, causing harmful outcomes, or introducing communication mistakes. Managers worried that outputs might drift from their “real” intent because tacit context cannot be cloned, with M10 explaining: “You can’t collect a lifetime of experience into a knowledge base, so some things the agent says won’t be what the leader truly wants”. Others recalled being held responsible for actions they had not authorized in their design fictions, as M6 noted: “You come back and find you’ve carried a lot of blame for things your agent said.” Workers, however, resisted being scapegoated. W2 asked, “If my AI made a decision that later caused problems, is the blame on me, on the AI, or on whom?” and W6 added, “If I revise work per the agent’s advice and it backfires, I’m probably the one who takes the fall.” Participants also anticipated “misunderstandings from transmission errors” (W1) and worried that agents might become a convenient excuse for poor human stewardship: “It gives people something to blame” (W3). As M8 summarized, “If they lean on my avatar and the output becomes distorted, and no one owns it, accountability becomes the issue.”
Participants also framed Manager Clone Agent use in terms of career opportunities, but from opposite vantage points. Some managers such as M1 and M6 imagined reinvesting time saved into managing up and outward: cultivating senior relationships, doing market visits, or expanding scope. Workers in the same workshops described the flip side: sparser access to managers, feeling devalued, and losing motivation. As W2 put it, “I kind of just don’t really feel like working anymore.” For workers, even routine updates or casual chats with supervisors were seen as crucial opportunities to demonstrate impact. W6 explained, “Weekly updates to the boss and some casual chat are themselves an opportunity to build rapport and leave a positive impression. You need to show your manager what impact your work has had, if you want a promotion.” Managers, meanwhile, feared that over-delegation to clone agents would cause skills to atrophy. As M2 warned, “People are greedy, the more they use it, the lazier they become.” To hedge this risk, others stressed the need to preserve core judgment for managers, as M8 insisted: “I want to keep control over my analysis of the metrics. AI can suggest targets, but I still set them.”
Participants further debated the extent to which managerial roles themselves might be replaced. Some workers argued that project managers with limited decision-making authority were highly vulnerable. W6 estimated “I think 80% of his work could really be replaced by an agent.” W2 similarly questioned the uniqueness of managerial labor, remarking: “I don’t think the division of labor made by AI or real managers will be any different, because I feel that neither is perfectly fair.”
Many managers such as M1, M3, M4, M6, M7, M10 and M11 defended higher-order decision-making, creative judgment, and relational work as difficult to automate. Yet some still admitted that agents might still perform adequately to some extent. As M9 reflected, “I would just have to resign to like it just wouldn’t be done maybe the way that I would do it. And that doesn’t mean it’s wrong. It can still do good enough.” Such reflections pushed managers to confront what M9 called “existential and ego connected identity questions,” echoed by M5: “I probably don’t want to authorize him because if my own work could be done by my agent, then what is this job position for me to do?”
7.2. Interpersonal Risks: Trust Cloned, But Wears with Distance
Trust in Manager Clone Agents initially built on trust in the human managers they represented. Participants explained that because the clone was seen as an extension of a known person rather than an anonymous system, the trust cultivated through prior human relationships could be partially transferred to the agent. In this sense, the “cloned” trust was borrowed credibility: workers felt more comfortable engaging with a Manager Clone Agent when they already trusted the manager behind it. Yet participants also emphasized that this form of trust was fragile. Unlike trust in a human being, which is reinforced by lived interactions, judgment, and emotional reciprocity, cloned trust rested on the assumption that the agent would faithfully mirror the manager’s intent and style. Once communication became mediated by agents, this assumption was tested and often strained by misinterpretations of value, uncertainty about authenticity, and the loss of affective and informal cues.
Participants first highlighted that the credibility of Manager Clone Agents depended on existing human relationships. M12 explained that “people will be more open with Manager Clone Agents than generic agents as it’s talking to someone you know vs. someone you don’t know.” Without this “underlying relationship,” M8 argued, the system would not constitute a “good functioning system.”
Participants then noted that trust weakened when managers relied too heavily on agents for direct interactions. They emphasized that agents could be useful to enhance communication, such as speeding up responses, filling gaps when managers were unavailable, or relaying straightforward information. But when small-group or one-on-one conversations were fully delegated to a clone, the meaning changed. Workers felt devalued, interpreting it as a signal that their work was “unimportant” (W2, W8) or that they were “unworthy of their manager’s time” (W7). This perception, M5 warned, could “diminish motivation.” W7 expressed the frustration more bluntly: “If you can’t be bothered to work with me at all, then why would I respect you?” Managers also worried about this dynamic in reverse, interpreting workers’ reliance on agents as avoidance or disengagement. As M4 explained in another perspective, “Managers wouldn’t like it if workers use a partition layer to block the imagined harm the boss might inflict on them.” In short, participants saw Manager Clone Agents as acceptable supplements to human communication but resisted their use as substitutes for genuine interaction.
Trust was further strained by uncertainty over authenticity. Workers such as W2 and W7 said they often could not tell whether a message came from their manager, from the agent itself, or whether their own contributions were being faithfully conveyed. Because Manager Clone Agents were designed to stand in for real people, this ambiguity directly undermined the very premise of the clone: workers needed to believe it faithfully mirrored their manager, yet could not be sure where the manager ended and the agent began. W10 illustrated this confusion on a storyboard (see Figure 2), showing the frustration that arises when overlapping identities blur accountability and when agents fail to recall past conversations between real individuals.
Figure 2. Story From W10 Showing Frustration When the Manager Clone Agent Fails to Recall Past Conversations.
Participants also emphasized the loss of emotional nuance and informal connection. W2 acknowledged that one benefit was the absence of negative outbursts, since an agent would not lose its temper. Yet even when agents simulated appropriate affect, workers often found it “unconvincing or even demotivating” (W1). The absence of genuine cues reduced moments of casual bonding. As W2 described, “At the beginning of meetings, there is usually a warm-up stage where we chat casually. I think this is when connections between people are built, and that gets lost interacting with agents.” This point sparked debate: M2 felt that authentic bonds were created outside formal settings such as meals or after-hours chats, while W2 considered meeting warm-ups essential. Managers also imagined longer-term erosion of interpersonal ties. M1 wrote in his design story a horrible relationship development that “The number of times I communicate with my staff has decreased, and to me their work is just a schedule sheet on a screen. I can’t even recall the names and appearances of newly hired employees anymore. The anonymous subordinate satisfaction feedback from HR is getting lower and lower.”
Participants stressed the need to balance efficiency with the preservation of interpersonal connection. As M8 remarked, “If I was leading a team, I’d want to make sure that personal connection was still there and that they don’t feel as if they’re just being offloaded.” In their view, trust that is cloned through Manager Clone Agents risks wearing away with distance, gradually reshaping the texture of workplace relationships and, eventually, organizational culture (see subsection 7.3).
7.3. Organizational Risks: Performance Up, Belonging Down
At the organizational level, participants described a paradox. They anticipated that Manager Clone Agents could improve systemic efficiency: reducing layers of coordination, redistributing managerial responsibilities, and broadening access to leadership. At the same time, they worried that these very efficiencies might erode the social fabric of organizations, weakening belonging and reshaping collective identity. In other words, while Manager Clone Agents promised to streamline structures, participants feared they might hollow out the cultural role that workplaces have long played.
Participants highlighted efficiency as a powerful driver of structural transformation. M10 described how Manager Clone Agents could take over repetitive approvals and routine tasks, relieving managers of “annoyances” and freeing up “an extra hour or two every day.” W6 imagined that if such delegation became common, “maybe there just aren’t that many managers anymore, the organization is becoming increasingly flat.” Echoing this, M10 noted that many intermediary roles such as HR staff or project managers exist mainly “to facilitate communication between managers and front line workers,” positions that agents could absorb. While some participants welcomed this flattening as a path to accessibility and speed, others feared the loss of “competition that historically drove innovation” (M6). Codifying rules into AI also raised new vulnerabilities. W7 warned in his design fiction that malicious workers could exploit system transparency, explaining, “I have now gotten away with embezzlement. I’ll move on to another company in about a year, because of tech turnover, and I’ll get to say I was trusted with a sizeable budget at a major AI-driven company.”
Beyond structure, participants emphasized the risks to organizational and community culture. W1 worried that when most communication flowed through Manager Clone Agents, “it greatly affect this sense of belonging.” M1 dramatized this erosion in a story where “dissatisfied employees were gradually replaced by those who no longer cared”. W2 pushed the scenario further, imagining a future where “The employees of the entire company are all working for AI,” leaving communal identity hollow. M5 satirized the endpoint: “An office filled entirely with mobile AI agents while humans stayed home: some enjoying more family time, others left isolated.”
8. Findings: Rethinking Manager Clone Agent Design in Future Workplace (RQ3)
Having surfaced both hopes and anxieties around Manager Clone Agents, participants turned toward imagining how such systems should be designed and deployed. Three interconnected design orientations emerged: keeping workers’ needs at the center (see subsection 8.1), positioning Manager Clone Agents as supportive extensions rather than replacements of managers (see subsection 8.2), and structuring agents along graduated levels of autonomy to allow for context-sensitive use and gradual trust-building (see subsection 8.3). These directions point to a future where Manager Clone Agents are not simply more efficient delegators, but carefully crafted collaborators whose roles are explicitly delimited to protect and nurture human relationships in teams and organizations.
8.1. Worker-Oriented: Prioritizing Workers’ Need
Workers repeatedly highlighted that the introduction of Manager Clone Agents does not only affect managers but also reshapes their own everyday realities. While managers may see such systems as convenient extensions of their authority, workers are the ones who must interact with them directly and continuously, who are managed by those clone agents, and who directly bear the consequences of the use of managerial clone agents. As M10 put it bluntly, most workers approach their jobs “just to earn a wage,” and if a digital agent suddenly appeared to “constantly command” them, making work “increasingly difficult” or adding “many annoying tasks,” they would “hate it deeply.” In such cases, rather than lightening workloads, the Manager Clone Agents multiplies expectations, leaving workers more burdened than before. Therefore, it’s highly important to take worker’s pains and needs into consideration when designing Manager Clone Agents.
Concerns about excessive control and surveillance were repeatedly emphasized. M11 observed that workers may feel as if they are reporting to “two leaders,” one being an extremely rational and emotionally detached version of their manager who “keeps asking endless questions.” M9, reflecting on her own workplace, stressed that she did not want colleagues to feel as if they were “always on the spot” or “watched the way that some companies just have your camera on all day.” For her, technology that fosters this feeling amounts to micromanagement.
Workers, positioned more vulnerably within workplace hierarchies, are especially attentive to whether Manager Clone Agents reinforce fairness and respect in their daily labor. W1, pointed to structural inequities that might emerge: “Workers willing to rely heavily on agents could offload routine labor and free up time for other pursuits, while those unwilling or unable to adopt the technology would be forced to shoulder more tasks themselves.” Over time, she argued, “such asymmetries could create divisions between groups of workers and amplify perceptions of unfairness.”
At the same time, participants highlighted that they care most about whether such systems deliver tangible value for workers. W3, for example, welcomed the idea that “An agent could improve and polish documents before sending them to the manager”. More broadly, as mentioned in section 6 workers expressed interest in Manager Clone Agents streamlining administrative tasks relaying feedback more consistently, etc. In these cases, the technology was seen as a supportive tool that reduced invisible labor.
In addition, whether Manager Clone Agents are perceived as supportive or extractive depends greatly on how their role is framed. The Manager Clone Agent’s role can be perceived differently as a tool or an assistant of the managers or the manager itself. As M12 summarized workplace dynamics often feel like “us versus them,” whether the Manager Clone Agents shows it cares about workers will shape how they respond to it.
These tensions further underscore the importance of clarity and consent. Workers repeatedly pointed out that the legitimacy of Manager Clone Agents depends on explicit agreements about where they may or may not be used. For example, W6, a junior programmer, argued that a one-on-one meeting is a crucial opportunity to “leave a good impression” and demonstrate initiative, and substituting an agent in this setting would feel “disrespectful.” W8 echoed this sentiment, insisting that in contexts “where emotional intelligence is important, I want the real person.” At the same time, he envisioned negotiating expectations directly with his manager: deciding together which interactions could be delegated to the agent and which must remain human-led.
Yet not all contexts lent themselves to easy consensus. M11, a Vice President of the company, described how he considered onboarding sessions fully replaceable: “It’s repetitive and a waste of time.” By contrast, W1, who’s only been working for 1-2 years, stressed that orientation was precisely when human presence mattered most: “If I’m a new employee joining a company, there might be many things that need to be coordinated with me, I would prefer a real person to handle that coordination with me.” Such disagreements illustrate how consent is not only about formal negotiation but also about uneven power and priorities. Managers may be motivated to offload repetitive duties, while workers may see those same moments as critical to relationship-building and trust. Because not all workers feel equally empowered to refuse a Manager Clone Agent’s presence, participants advocated for organizational safeguards, such as granting workers the protected right to request “human-only” interactions or limiting the contexts in which Manager Clone Agents may operate.
8.2. Revisit “Enhance vs. Replace”: Manager Clone Agents as Interpersonal Bond
The HCI community has long advocated for AI that enhances rather than replaces human capabilities (Capel and Brereton, 2023). Our findings extend this debate by showing how the line between enhancement and replacement is inseparable from the quality of manager–worker relationships. Participants judged the value of Manager Clone Agents less by efficiency alone than by whether these systems strengthened or weakened the bonds between humans. Importantly, we found sometimes managers and workers often imagined enhancement in different terms, reflecting their divergent stakes in organizational life.
For managers, enhancement meant receiving support that extended their capacity without threatening authority or identity. M9 expressed her wish for the agent to help her “be a better manager,” for instance by reminding her when she was “over budget” or helping her track priorities, rather than taking over leadership itself. Similarly, M11 stressed that the agent should remain in a service role, “approving a travel request if a boss is unavailable,” but not as “always pointing fingers.” M8 echoed this distinction, noting she was comfortable when the system was “clarifying what’s already there,” but grew uneasy when it began “creating new responses,” since that crossed into authorship and identity.
Workers, by contrast, emphasized enhancement as a way to maintain access and grow. W3 imagined using a Manager Clone Agent to manage emotions before approaching her boss: when she feared she might “lose control” in a heated moment, the agent could help frame her message more constructively. W8 described the agent as “an extension of the manager” that could provide information or quick decisions while still leaving escalation to the human manager. For workers, enhancement thus meant greater availability and emotional scaffolding, rather than authority protection.
Yet the line between enhancement and replacement remained blurry. W4’s story illustrated this dilemma: “I asked, ‘Then how should I communicate this?’ The Agent replied, ‘Do you want me to pass it on or do you want to do it yourself?’ I said, ‘What do you think?’ The Agent said, ‘If you do it, it can help you practice your communication skills, but you might not get what you want. But if I do it for you, it might be more effective but you wouldn’t be involved in the process.’” While workers saw this as a trade-off between empowerment and convenience, managers like M3 worried that such mediation of clone agents could “build an imaginary wall that ends up hurting both sides,” deepening misunderstanding rather than reducing it.
Even as participants offered these visions of enhancement, they consistently stressed the importance of retaining human autonomy. M8 argued that preserving creativity and trust within teams is critical, noting that relationships, respect, and psychological safety “take time to develop” and cannot be delegated to machines. She warned that if technology undermines “any version of I can’t trust it, or I don’t know, or I don’t feel comfortable,” it risks eroding the interpersonal foundations that drive productivity. M9 echoed this sentiment by emphasizing the value of choice. She wanted to “see it all, but be able to say like, I’m not busy right now, but I trust that bot to do the answering for me or I’ll get back to you in 5 minutes.”
8.3. Tiered Autonomy: Gradual Trust-Building and Flexible Configuration
As participants imagined how Manager Clone Agents might operate in practice, they often turned to the metaphor of “levels of autonomy” familiar from self-driving cars (M2, M10). Rather than treating Manager Clone Agents as a single, fixed entity, they described autonomy as a spectrum, structured along dimensions such as representation, proactivity, and delegation. These dimensions offered a shared language for debating what forms of cloning would feel acceptable, what boundaries should be respected, and how much authority agents ought to exercise in different workplace contexts.
8.3.1. Representation.
Participants emphasized the spectrum of representational fidelity, ranging from highly stylized, cartoon-like avatars to hyper-realistic deepfakes indistinguishable from the human manager. For modality, At the lowest level, Manager Clone Agents may exist as text-based entities; at higher levels, they may incorporate voice, video, or even embodied presence in virtual or physical environments. While M13 expressed discomfort with Manager Clone Agents that “perfectly replicated their speech patterns and idiosyncrasies”, finding the imitation “uncanny and inauthentic”, workers similarly noted heightened anxiety when they could not discern whether communication originated from the manager or the agent (W2, W7). M3, W7 and W8 agreed that abstract forms of representation, such as “cartoon styled (M3) or “orb-like” voice interface (W7, W8), were perceived as less deceptive and more broadly acceptable. W8 further explained that embodiment affects comfort: “If it’s an avatar, it requires more attention, more effort to be present, like I’m performing extra socially.” Text-based agents, by contrast, were viewed as lightweight and flexible, more easily integrated into everyday practices such as Slack channels.
8.3.2. Proactivity.
Another dimension concerns the level of initiative that Manager Clone Agents should exercise, which participants saw as varying by task. At the lowest level, agents might simply observe or log interactions without intervening. At intermediate levels, they could relay pre-authorized statements or provide clarifications. The highest levels would involve spontaneous contributions or even steering discussions that many participants found unacceptable for sensitive or high-stakes contexts. As noted in subsection 8.2, M8 and M9 worried about being misrepresented by agents making statements they had not endorsed, or by discussions diverted in their absence in ways that later shaped decisions. M10 cautioned that disclosure must be carefully limited, warning, “The authority to disclose information should actually be tightened as well. For example, if it involves company finances and someone says that cash flow is tight, workers might panic.” Similarly, M11 argued that strong presence could backfire sometimes, suggesting, “Don’t make their presence too strong. Even just occasional reminders or emotional support are enough to help everyone gradually accept this new colleague. If their presence is too strong, everyone will feel threatened.”
8.3.3. Delegation.
Finally, the scope of action delegated to Manager Clone Agents emerged as a central concern. Low-delegation configurations would confine agents to communication support, while high-delegation systems could independently execute managerial tasks, such as reallocating resources or authorizing workflows, without human intervention. M7 stressed limits: “As a strictly constrained ‘assistant,’ emphasize ‘reminders’ rather than ‘decisions,’ with transparency and explainability prioritized.” M10 described the risks of overstepping: “Some overly specific suggestions should be constrained. The remaining decision space should be left to the person doing the work. Otherwise, it’s like having a CTO constantly coaching right next to the coder, which makes workers uncomfortable.” M5 proposed explicit labeling: “Label the level of an authorized agent, for example, before getting a Gold level no one pays attention, but after getting Gold any command can be executed. So you can design agent tiers to let the interlocutor understand their current permissions.”
Collectively, participants agreed that no single configuration of autonomy is universally appropriate in all teams and organizations. Instead, Manager Clone Agent design must accommodate contextual variation across organizations, teams, and leadership styles. Just as workplace leadership itself is heterogeneous, Manager Clone Agents should preserve space for managers to calibrate their use in ways that best fit their teams. Importantly, dimensions of autonomy do not operate in isolation but in combination. For example, “a highly realistic Manager Clone Agent that only listens or relays pre-approved messages may still be experienced as low-autonomy” (M9), while “a less anthropomorphic text-based agent that intervenes proactively may feel disproportionately intrusive” (W8). Fine-tuning these combinations is therefore essential for aligning Manager Clone Agent design with individual comfort and organizational needs, while supporting gradual trust-building and long-term adoption.
9. Discussion
This study examined how managers and workers envisioned the possibilities and risks of Manager Clone Agents in future workplaces. Drawing on workshop stories and discussions, we found that participants situated Manager Clone Agents at the intersection of productivity and relations: managers saw them as extensions of their authority and capacity, while workers emphasized how these systems would reshape recognition, fairness, and everyday labor. Across individual, interpersonal, and organizational levels, participants oscillated between hopes of enhanced productivity and fears of displacement or relationships erosion. Building on these findings, we contribute to ongoing debates in HCI and organizational research by situating Manager Clone Agents in three broader conversations: the distinctiveness of clone agents from other algorithmic management systems (see subsection 9.1), the rethinking of managerial work in the face of automation (see subsection 9.2), and the reconfiguration of worker–manager relations in an automated workplace (see subsection 9.3).
9.1. From Algorithmic Bosses to Cloned Managers: Reframing Managerial Authority in AI Systems
Much of the HCI and organizational literature on algorithmic management has focused on so-called algorithmic bosses, systems that allocate tasks, monitor workers, and enforce compliance in platform labor and other digitally mediated settings (Möhlmann et al., 2021; Jarrahi et al., 2020; Lee, 2016). These systems are typically characterized by impersonality: they operate through metrics, dashboards, and automated rules, with limited reference to the identities of specific human managers. Trust and legitimacy in such systems are often evaluated in terms of procedural fairness, transparency of rules, or predictability of outcomes (Lee et al., 2019b; Prassl, 2019).
Our findings extend this line of work by showing how Manager Clone Agents powered by GenAI, as a new form of AI management, are being re-imagined in fundamentally different terms. Unlike faceless algorithms, Manager Clone Agents are not abstract authorities but personalized proxies that explicitly embody a named individual. They carry the symbolic presence, voice, and persona of a particular manager, and in doing so transform how participants judged their legitimacy. For workers, what mattered was not only whether the system allocated tasks effectively and fairly, but whether it felt like an authentic extension of their manager. This personalization distinguishes Manager Clone Agents from prior models of algorithmic authority.
At the same time, the same symbolic link that granted Manager Clone Agents legitimacy could also undermine it. Both workers and managers expressed concern about the risks of misrepresentation, where an agent might speak or act in ways the human manager did not intend, thereby creating confusion over accountability. Participants stressed that while they valued clarification or reinforcement of existing managerial decisions, they were uneasy when Manager Clone Agents generated new responses or independent contributions, as this crossed into authorship that challenged identity and authority. These concerns highlight a paradox: the closer Manager Clone Agents come to cloning the manager, the higher the stakes when authenticity is uncertain.
This distinction suggests that Manager Clone Agents occupy a different conceptual space from algorithmic bosses or generic virtual assistants. Rather than replacing managerial oversight with impersonal rules, they aim to reproduce the relational authority of a specific human. This raises new questions for HCI and organizational theory about how authority is mediated, shared, or contested when a system is designed not simply to manage work, but to embody the manager. Our contribution here is to show that the design of AI systems in organizational contexts cannot be understood solely through the lens of efficiency, fairness, or control. With Manager Clone Agents, legitimacy and trust hinge on whether the technology can sustain the symbolic and relational bonds that management entails.
9.2. Rethinking Managerial Work in the Context of Automation
Research on managerial automation has often assumed a clear division between routine tasks that can be automated and higher-level strategic or relational work that should remain human-led (Shao et al., 2025; Zaidi et al., 2025; West, 2018). This division frames automation as a way of freeing up managers to focus on strategy, creativity, or leadership by removing administrative burdens (Raisch and Krakowski, 2021). Our findings complicate this assumption by showing that even routine managerial tasks, such as on-boarding new employees or providing quick feedback, carry symbolic and relational weight for workers. When delegated to Manager Clone Agents, they risk being reinterpreted as a lack of personal engagement, thereby weakening the interpersonal bonds that underpin managerial legitimacy.
Participants such as W3, W4, W8, M5 and M10 also imagined Manager Clone Agents as amplifiers of managerial capacity, as we showed in subsection 6.4, but again this was framed through relationships. By handling repetitive approvals or relaying pre-defined information, Manager Clone Agents were seen as allowing managers to invest more in external relationships and long-term vision. Yet participants insisted that the legitimacy of this amplification depended on boundaries. They stressed that workers looked to managers not only for efficiency, but also for discretion, authorship, and relational assurance that decisions came from a trusted human source. Delegating too far into the relational domain was perceived as hollowing out the leadership identity and undermining interpersonal trust between managers and teams.
These insights call for a rethinking of managerial work in the rise of GenAI. Management cannot be reduced to a bundle of discrete tasks, neatly divided into automatable and non-automatable categories. Instead, our findings emphasize its fundamentally relational character (Sluss and Ashforth, 2007). Managerial labor involves decision-making, but also symbolic and affective practices that maintain respect, recognition, and psychological safety within teams. For HCI, this shifts the challenge from designing automation that maximizes efficiency to designing systems that preserve the relational dimensions of leadership. By surfacing these nuances, our study contributes to a richer understanding of what aspects of management can be augmented by Manager Clone Agents and what must remain human to sustain the interpersonal bonds that give managerial work its meaning.
9.3. Rethinking Worker–Manager Relations in the Age of AI Agents
Another contribution of our study is to revisit how automation reshapes the relational fabric between workers and managers. Research on algorithmic management has often conceptualized this relationship as one of asymmetrical control, where digital systems strengthen managerial oversight while limiting worker agency (Barati and Ansari, 2022; Zhang et al., 2025; Xiao et al., 2025a). Our findings complicate this view by showing that Manager Clone Agents are imagined by our participants not only as tools of authority, but also as relational intermediaries that can redistribute opportunities, recognition, and burdens in uneven ways.
For the managers in our participants pool, Manager Clone Agents were valued as extensions of authority and capacity. They promised responsiveness in the manager’s absence, relief from repetitive administrative demands, and the ability to scale guidance to larger teams. Workers, however, approached Manager Clone Agents from a different vantage point: what mattered to them was whether the system preserved access to managers, created fair opportunities for visibility, and supported their career growth. While managers worried about identity and authorship, workers worried about recognition, legitimacy, and respect. These diverging stakes underline that automation in management cannot be understood solely through the lens of leadership efficiency but must account for its downstream effects on workers’ sense of belonging and advancement.
Our study contributes to a richer theorization of worker–manager relations in the age of AI agents. Rather than assuming automation uniformly empowers managers or uniformly disempowers workers, Manager Clone Agents highlight how authority, recognition, and trust are continually renegotiated in mediated interactions. For HCI and organizational research, this suggests that the legitimacy of AI in management hinges not only on technical reliability but on its ability to sustain relational equity between workers and managers.
9.4. Design and Policy Implications
Beyond theoretical contributions, our findings also highlight several implications for the design and governance of Manager Clone Agents. From a design perspective, Manager Clone Agents should incorporate mechanisms that clarify boundaries and preserve human agency. Participants valued augmentation but resisted substitution, suggesting that systems should make it explicit when an interaction is mediated by an agent, what kinds of contributions the agent is authorized to make, and how accountability is distributed. Design strategies could include graduated levels of initiative, ranging from lightweight reminders to bounded forms of delegation, so that managers and workers can calibrate the degree of automation appropriate to specific contexts. Equally important are features that enable workers to opt out, request human-only interactions, or signal when relational presence is essential, such as in mentoring or evaluation.
From a policy perspective, organizations adopting Manager Clone Agents must create safeguards that protect workers from new vulnerabilities. This includes ensuring that workers’ reliance on or resistance to Manager Clone Agents does not become a hidden basis for performance evaluation, and that managerial delegation does not erode opportunities for worker recognition and advancement. Policies should also mandate transparency about when and how Manager Clone Agents are deployed, as well as clear escalation paths for contesting errors or misrepresentations. At a broader level, regulatory frameworks may need to address the dual identity of Manager Clone Agents as both tools and proxies, clarifying where legal and ethical responsibility lies when agents act in a manager’s name.
10. Limitations and Future Work
Like any qualitative and speculative study, our work has several limitations that should be acknowledged. First, our findings are based on six design fiction workshops with 23 participants. While this approach enabled us to surface rich imaginaries and relational concerns, the sample is not representative of all industries, organizational sizes, or cultural contexts. Second, our use of speculative design elicited visions of possibilities and risks rather than documenting actual deployments. This allowed participants to reflect openly without being constrained by current technological limits, but it also means that their narratives may differ from how Manager Clone Agents are eventually integrated in practice. Third, our focus was primarily on managers and workers situated within formal organizations. Other stakeholders, such as HR professionals, policymakers, or AI developers, are likely to shape how Manager Clone Agents are designed and governed. Expanding the scope to include these perspectives would enrich understanding of the broader ecosystem in which Manager Clone Agents operate. at they enhance rather than erode the relational and cultural foundations of work.
11. Conclusion
This paper examined how managers and workers envisioned the opportunities and risks of Manager Clone Agents, AI-powered surrogates trained on managerial communication and practices to act on behalf of leaders. Through six design fiction workshops with 23 participants, we identified potential supportive roles in future workplaces and the tensions they created across individual, interpersonal, and organizational levels. We argue that the design and governance of Manager Clone Agents must be grounded in relational awareness, ensuring that these systems enhance rather than replace the symbolic, affective, and interpersonal practices that give managerial work meaning.